OpenStudy (eric_d):

If y=x^2e^x, find dy/dx and show that dy/dx=0 when x=0 or x=-2.

jimthompson5910 (jim_thompson5910):

use the product rule to find dy/dx

OpenStudy (eric_d):

d/dx uv = u d/dx v + v d/dx u ?

OpenStudy (eric_d):

Apply the above formula ?

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (eric_d):

So, u=x^2 and v= e^x

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (eric_d):

 like this ??